OpenStudy (anonymous):

How do you Find the vertex, focus, and directrix of the parabola and how do you sketch its graph? x + y^2 = 0

OpenStudy (anonymous):

start by writing \(-y^2=x\)

OpenStudy (anonymous):

hmm actually i think you might want the general form as \((y-k)^2=4p(x-h)\) no matter because the vertex is still \((0,0)\)

OpenStudy (anonymous):

the reason to write it as \((y-k)^2=4p(x-h)\) is that \(p\) tells you the focus. in this case we have \(y^2=-x\) and so \(4p=-1\) making \(p=-\frac{1}{4}\) so your focus is \((-\frac{1}{4},0)\) i.e. \(\frac{1}{4}\) units to the left of the vertex

OpenStudy (anonymous):

sorry i screwed up on first post. it should have been \((y-k)^2=4p(x-h)\) from the start and your job is to force your equation to look like that, i.e. interperet \[y^2=-x\] in terms of \[(y-k)^2=4p(x-h)\]

OpenStudy (anonymous):

since the focus is at \((-\frac{1}{4},0)\) to the right of the vertex, we know the directrix is on the other side, the same distance away, so it is the line \(x=\frac{1}{4}\)

OpenStudy (anonymous):

we can see it from the nice picture http://www.wolframalpha.com/input/?i=parabola+y^2%3D-x

OpenStudy (anonymous):

wow, another typo, \((-\frac{1}{4},0)\) is to the LEFT over the vertex, so the directrix is to the right. correct answer though